My son recently took a test an had the following question wrong:
3x^2 - 4 = 0 That's 3 "x" squared. He was require to provide the solution and the correct answer is: +/- 2 times the square root of 3 divided by 3
How does one come upon that answer? He is having difficulty trying to find out how to begin

Answer by stanbon(75887)

3x^2 - 4 = 0
---
3x^2 = 4
--
x^2 = 4/3
---
x = +/-(2/sqrt(3))
---
Rationalizing the denominator you get:
x = +/- (2sqrt(3))/3
---
OR you could write it as:
x = (2/3)sqrt(3) or x = (-2/3)sqrt(3)
==========================================
Cheers,
Stan H.
============